suppose the zero of the polynomial function f(x) are -2,0,3 what can be the equation of f(x)
Answer #1

x=0 n = x=0 The zeroes are - 2 x=3 (71+2) =0 171-3)=0 2o the factors are (2+2), (4), (1-3) If we multiply these factors we ge